# FLawN-T5: An Empirical Examination of Effective Instruction-Tuning Data Mixtures for Legal Reasoning

### v1

@@ Abstract
Instruction tuning is an important step in making language models useful for direct user interaction. However, many legal tasks remain out of reach for most open LLMs and there do not yet exist any large scale instruction datasets for the domain. This critically limits research in this application area. In this work, we curate LawInstruct, a large legal instruction dataset, covering 17 jurisdictions, 24 languages and a total of 12M examples. We present evidence that domain-specific pretraining and instruction tuning improve performance on LegalBench, including improving Flan-T5 XL by 8 points or 16% over the baseline. However, the effect does not generalize across all tasks, training regimes, model sizes, and other factors. LawInstruct is a resource for accelerating the development of models with stronger information processing and decision making capabilities in the legal domain.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) advanced significantly, evident in their performance gains across numerous benchmarks, including SuperGLUE (Wang et al., 2019), MMLU (Hendrycks et al., 2021b), and various human examinations (OpenAI, 2023), such as the U.S. bar exams for law practice admission (Katz et al., 2023). However, the interplay between domain-specific training and within-domain evaluation is poorly understood. This work examines how training on domain-specific legal corpora affects performance on the widest set of legal-domain evaluation benchmarks known to the authors. We thus conduct a study of the ability of models to answer questions, classify, make judgments, extract information, and otherwise perform decision making or higher-order cognitive tasks (i.e., to “reason”) within a limited domain, as opposed to broad-domain benchmarking. We present evidence that domain-specific pretraining and instruction tuning improve performance—but the effect does not generalize across all tasks, training regimes, model sizes, and other factors.
@@
To study this, we use the MultiLegalPile (Niklaus et al., 2023b), a 689GB multilingual legal corpus, for continued pretraining. Because no instruction dataset for legal reasoning is available, we introduce LawInstruct, spanning 24 languages in 17 jurisdictions on four continents. It contains 12M training examples for QA, entailment, summarization, and information extraction tasks in the legal domain, each presented as a bespoke instruction with corresponding output. With this large instruction dataset in hand, we fine-tune models and then perform quantitative analyses of their outputs on the LegalBench (Guha et al., 2023) and MMLU (Hendrycks et al., 2021a) benchmark suites. Instruction tuning Flan-T5 models on LawInstruct, we achieve a balanced accuracy of 58.1 on LegalBench for the XL size, improving by 8 points or 16% over the baseline. The Small model even improves by 9.6 points or 38.1% and by 14 points or 55.4% when we also continue pretraining it.
The contributions of this paper are four-fold: First, we curate the first legal instruction dataset by unifying and writing instructions for 58 high-quality annotated datasets covering diverse legal tasks. Second, we continue pretraining and instruction tune T5, mT5, and Flan-T5 models and achieve new state-of-the-art on LegalBench in all tested parameter ranges. Third, we perform a wide range of ablations across different dataset configurations providing new insights for domain adaptation. Finally, we publicly release the permissively-licensed portion of the curated dataset on the Hugging Face Hub [^1] https://huggingface.co/lawinstruct/lawinstruct and release the code used to create the dataset [^2] https://github.com/JoelNiklaus/LawInstruct/ including pointers on how to access the portions of the data that require special agreements.
@@ 2 Experimental Setup
In this section, we describe the experimental setup we used to test the effect of pretraining and instruction tuning on in-domain legal data. We consider T5 v1.1+LM adaptation (Raffel et al., 2020; Lester et al., 2021), Flan-T5 (Chung et al., 2022) and mT5 (Xue et al., 2021) models in the sizes Small, Base, XL and XXL, allowing us to study effects over different model scales. We selected the T5 family of models over other models for three reasons: 1) Flan-T5 XL and XXL perform best in their parameter range on LegalBench, 2) T5 and mT5 allow us to measure the effect of multilinguality in a controlled setting, and 3) the T5 model family contains models from 60M parameters (Small) to 11B (XXL) allowing us to study scaling behaviour also at smaller scales.
We used random seed 42 throughout. Our experiments were performed with T5X  [^3] https://github.com/google-research/t5x on TPUv4 pods using between 2 and 512 cores. We present the mean across tasks per LegalBench category and for LegalBench overall by aggregating over the categories.
@@ 2.1 Continued Pretraining
We continue pretraining on the MultiLegalPile (Niklaus et al., 2023b), a 689GB corpus in 24 languages from 17 jurisdictions. The MultiLegalPile corpus, which includes diverse legal data sources with varying licenses, allows for pretraining NLP models under fair use, with more permissive licenses for the Eurlex Resources and Legal mC4 subsets. It consists of four large subsets: a) Native Multi Legal Pile (112 GB), b) Eurlex Resources (179 GB), c) Legal mC4 (106 GB), and d) Pile of Law (292 GB). For our experiments on the mT5 models, we use the entire corpus, and for the experiments on the T5 and Flan-T5 models we use only the texts in the English language.
We continued pretraining (a.k.a. domain adaptation of) the models with 512 tokens in both inputs and targets on the MultiLegalPile (Niklaus et al., 2023b) whereas the original models were pretrained on C4 (Raffel et al., 2020). We used the UL2 mixture (Tay et al., 2022) due to its promise to enable improved training efficiency with its mixture of denoisers tasks. In initial experiments we used batch size 1024 and warmed up the learning rate linearly for the first 10K steps from 2.5e-3 to 5e-3 and finally decayed down to 1.5e-3. However, we noticed training instabilities for the XXL models. We switched to a constant learning rate of 1e-3 and ran a sweep over batch sizes 64, 128, 256, 512, 1024. The XXL model only trained stably with batch size 128.
@@ 2.3 Evaluation
We evaluate our models on LegalBench and MMLU to test in-domain and generalization performance, respectively. LegalBench (Guha et al., 2023) consists of 162 tasks evaluating different aspects of legal classification and reasoning. Each task is assigned to one of six categories, depending on the broader type of legal reasoning implicated. LegalBench tasks are sourced from both previously constructed datasets and novel tasks collected from different members of the legal community (e.g., lawyers, legal impact organizations, legal academics). As such, LegalBench is thought to capture tasks of interest and practical applicability. LegalBench tasks span a wide range of legal subject areas (e.g., contracts, civil procedure, tax, etc.) and text-types (natural language, contractual terms, judicial opinions, etc.). The majority of tasks are either classification or extraction tasks, thus enabling automated evaluation. Massively Multilingual Language Understanding (MMLU) benchmarks models factual knowledge (Hendrycks et al., 2021a). MMLU contains multiple-choice questions on 57 subjects, including three related to law: jurisprudence, international law, and professional law. While there exist multilingual benchmarks like LEXTREME (Niklaus et al., 2023a), they are still very challenging for generative models not specifically finetuned per task. Therefore, we focus on LegalBench and MMLU both in the English language.
For evaluation, we set temperature to 0 in line with accepted practice for LegalBench evaluation (Guha et al., 2023) that focuses on the highest-likelihood token sequence with minimal variance. We removed the following prefixes before scoring: “label”, “target”, “option”, “answer”, “a:”. We did not evaluate Rule QA because it necessitated manual evaluation. We show paper baseline results compared with our runs in Appendix D Table 3. Our XL model is quite close to the XL model in the LegalBench paper, but there are significant differences for the XXL model. We provide a more detailed analysis of possible causes in Section B.1. Unless specifically mentioned, we compare to our baselines results. We hold out LegalBench tasks overlapping with LawInstruct tasks unless specified otherwise (see Section B.2 for details).
@@ 3 Results
In this section, we discuss the main results achieved by instruction tuning and continued pretraining.
@@
Figure 5: Performance progression on LegalBench and MMLU from baseline to instruction finetuning and domain adaptation + instruction finetuning.
Table 1: Progression of performance from baseline to instruction tuning (IFT) and instruction tuning together with continued pretraining (PRE-IFT+IFT).
@@
To analyze the change in performance in more detail, we show the difference to the baseline for the XL model on LegalBench and MMLU across tasks (see Figure 6) and across categories (see Figure 7). We find that FLawN-T5 outperforms baseline Flan-T5 in most LegalBench tasks in most categories. The exception are tasks in the interpretation category, specifically CUAD (Hendrycks et al., 2021c), where the fine-tuned model is actually worse than the baseline by around 10 points on average. A possible explanation could be negative transfer from the instruction tuning data since the task formulations are very different to the instructions in LegalBench. In MAUD (Wang et al., 2023a) and Contract-NLI (Koreeda & Manning, 2021), the instructions are much more similar from LawInstruct to LegalBench, leading to improvements compared to the baseline. On MMLU, most categories and tasks see increases in performance, especially the categories social sciences and other. We find that performance suffers mostly in the STEM category and to some extent in the humanities. Interestingly, the largest drop is in machine learning but the largest rise is in high school computer science. In the humanities, more “hard” disciplines are affected by performance decrease, such as formal logic and logical fallacies.
@@ 4 Ablations
In this section, we present diverse ablations across the starting checkpoints, data mixtures, sampling styles, licences, instruction styles and amount of instruction tuning data during pretraining. We show additional ablations regarding crosslingual transfer from multilingual data in Appendix C.
@@ 4.1 Starting Checkpoint
Should you start in-domain instruction tuning from a base model or from an instruction tuned model? $\Rightarrow$ Starting from an instruction tuned model is better across sizes except Small. In Figure 8, we compare instruction tuning from a base T5 and a Flan-T5 model in four different sizes (Small, Base, XL and XXL) (detailed results in Appendix D Table 4). We find that for the larger sizes, the instruction tuned Flan-T5 is a better starting point ($p<0.001$), leading to higher performance on LegalBench. For the Small size the difference is not statistically significant ($p=0.058$). By default, we use the Flan-T5 model as a starting point in all following experiments unless specified otherwise.
@@ 4.2 Data Mixture
What data mixtures should you choose for in-domain instruction tuning? $\Rightarrow$ Mixing in general instruction tuning datasets is necessary. In Figure 9, we compare instruction tuning with three different data mixtures: lawinstruct, flan2 (Chung et al., 2022), and flan2-lawinstruct (where we sample equally from flan2 and lawinstruct) (detailed results in Appendix D Table 5). Interestingly, when only training on lawinstruct, downstream accuracy drops, possibly due to the instructions in our datasets being formulated differently than the original Flan instructions. Training on flan2 and flan2-lawinstruct leads to an aggregate increase of 7.7 points (48.3 to 56) and 10.8 points (48.3 to 59.1) respectively. By default we use the flan2-lawinstruct mixture in all following experiments unless specified otherwise.
4.3 Sampling Style
Should we sample each dataset equally or rather by the number of examples? $\Rightarrow$ Sampling by the number of examples generally leads to better performance. In Figure 10, we compare the performance of two sampling styles (equal sampling of each dataset and sampling by the number of examples) across both the research and commercial licensed dataset (detailed results in Appendix D Table 6). For the XL and XXL sizes, sampling by the number of examples is better than equal weight for datasets for both the research and commercial datasets, although not always statistically significant (XL res $p=0.049$, XL comm $p=0.052$, XXL res $p<0.001$, XXL comm $p=0.31$). For the Small size, sampling by the number of examples is better for the research dataset ($p<0.001$) but not for the commercial dataset ($p=0.099$), while there is no difference for the Base size. By default, we sample by the number of examples in all following experiments unless specified otherwise.
4.4 License of Instruction Tuning Datasets
Do we need data licensed non-commercially for good performance? $\Rightarrow$ The commercially licensed data seems to be enough for the larger models. In Figure 10, we compare the performance of two differently licensed datasets (research and commercial licenses) across both sampling each dataset equally and by the number of examples (detailed results in Appendix D Table 6). Except for Small size ($p<0.001$), using more diverse data available only for research shows no significant benefit. By default, we use the commercially licensed dataset in all subsequent experiments unless specified.
@@ 4.5 Instruction Style
Are models trained with more diverse instructions better on LegalBench? $\Rightarrow$ Results are mixed, overall just using one instruction is probably sufficient. In Figure 11, we compare the performance of training with just one manually written instruction vs. ten paraphrased instructions with GPT-4 from one seed instruction (detailed results in Appendix D Table 8). For Flan-T5 (see Table 8), for Small, one instruction is better than ten ($p=0.035$); for the other sizes we find no difference. For mT5 (see Figure 11(b)), for Small, one instruction is worse than ten both monolingual ($p=0.005$) and multilingual ($p=0.01$) whereas for XL, ten English instructions underperform one English ($p<0.001$) and ten multilingual ones ($p<0.001$). In aggregate, differences are small without a consistent trend.
@@ 4.6 Amount of Instruction Data During Continued Pretraining
How much instruction tuning data should be mixed in during continued pretraining? $\Rightarrow$ Continued pretraining seems to be rather robust with regard to the amount of instruction tuning samples mixed in. In Tables 10, 11, 12 and 13, we investigate the benefit of mixing varying amounts of instruction tuning data in during continued pretraining (detailed results in Appendix D Tables 10, 11, 12 and 13). We compare results on LegalBench of instruction tuning runs after 10K to 90K steps of continued pretraining. For the Small model, the benefit of continued pretraining over just instruction tuning is significant (34.9 for just instruction tuning vs.  40 after continued pretraining). Conversely, for the XL model, continued pretraining often underperforms compared to just instruction tuning. For the XXL model, more instruction tuning samples during continued pretraining improve performance, unlike for the Small and XL models. Across sizes, continued pretraining’s effectiveness appears robust to the number of instruction tuning samples used. [^4] Just mixing instruction tuning data during continued pretraining without more instruction tuning does not improve results.

### v2

# LawInstruct: A Resource for Studying Language Model Adaptation to the Legal Domain

@@ Abstract
Instruction tuning is an important step in making language models useful for direct user interaction. However, the legal domain is underrepresented in typical instruction datasets (e.g., only 10 out of 1600+ tasks in Super-NaturalInstructions). To study whether instruction tuning on legal datasets is necessary for strong legal reasoning, we aggregate 58 annotated legal datasets and write instructions for each, creating LawInstruct. LawInstruct covers 17 global jurisdictions, 24 languages and a total of 12M examples across diverse tasks such as legal QA, summarization of court cases, and legal argument mining. We evaluate our models on LegalBench, measuring legal reasoning across five categories in 162 challenging and realistic legal tasks, and MMLU, to measure potential drops in general reasoning capabilities. We find that legal-specific instruction tuning on Flan-T5 – yielding FLawN-T5 – improves performance on LegalBench across all model sizes, with an aggregate increase of 15 points or 50% over Flan-T5 for the base size. No model size shows performance drops in MMLU. We publish LawInstruct as a resource for further study of instruction tuning in the legal domain.
LawInstruct: A Resource for Studying Language Model Adaptation
to the Legal Domain
Joel Niklaus ${}^{S\,X}$   Lucia Zheng S   Arya D. McCarthy J   Christopher Hahn X   Brian M. Rosen X
[…]
@@ 1 Introduction
In recent years, Large Language Models (LLMs) advanced significantly, evident in their performance gains across numerous benchmarks, including SuperGLUE (Wang et al., 2019), MMLU (Hendrycks et al., 2021a), and various human examinations (OpenAI, 2023), such as the U.S. bar exams for law practice admission (Katz et al., 2023). However, the interplay between domain-specific training and within-domain evaluation is poorly understood. This work examines how training on domain-specific legal corpora affects performance on the widest set of legal-domain evaluation benchmarks known to the authors. We thus conduct a study of the ability of models to answer questions, classify, make judgments, extract information, and otherwise perform decision making or higher-order cognitive tasks (i.e., to “reason”) within a limited domain, as opposed to broad-domain benchmarking. We present evidence that domain-specific pretraining and instruction tuning improve performance—but the effect does not generalize across all tasks, training regimes, model sizes, and other factors.
@@
To study this, we use the MultiLegalPile (Niklaus et al., 2024), a 689GB multilingual legal corpus, for continued pretraining. Because no instruction dataset for legal reasoning is available, we introduce LawInstruct, spanning 24 languages in 17 jurisdictions on four continents. It contains 12M training examples for QA, entailment, summarization, and information extraction tasks in the legal domain, each presented as a bespoke instruction with corresponding output. With this large instruction dataset in hand, we fine-tune models and then perform quantitative analyses of their outputs on the LegalBench (Guha et al., 2023) and MMLU (Hendrycks et al., 2021b) benchmark suites. Instruction tuning Flan-T5 models on LawInstruct, we achieve a balanced accuracy of 58.1 on LegalBench for the XL size, improving by 8 points or 16% over the baseline. The Small model even improves by 9.6 points or 38.1% and by 14 points or 55.4% when we also continue pretraining it.
The contributions of this paper are four-fold: First, we curate the first legal instruction dataset by standardizing and writing instructions for 58 high-quality annotated datasets covering diverse legal tasks to make them usable for instruction tuning in the first place. Second, we continue pretraining and instruction tune T5, mT5, and Flan-T5 models and achieve new state-of-the-art on LegalBench in all tested parameter ranges. Third, we perform a wide range of ablations across different dataset configurations deepening our understanding of adapting models to specific domains. Finally, we publicly release the permissively-licensed portion of the curated dataset on the Hugging Face Hub [^1] https://huggingface.co/lawinstruct/lawinstruct and release the code used to create the dataset [^2] https://github.com/JoelNiklaus/LawInstruct/ including pointers on how to access the portions of the data that require special agreements.
@@ 2 Experimental Setup
In this section, we describe the experimental setup we used to test the effect of pretraining and instruction tuning on in-domain legal data. We use random seed 42 throughout. Our experiments were performed with T5X  [^3] https://github.com/google-research/t5x on TPUv4 pods using 2 to 512 cores. We present the mean across tasks per LegalBench category and for LegalBench overall by aggregating over the categories. We consider T5 v1.1+LM adaptation (Raffel et al., 2020; Lester et al., 2021), Flan-T5 (Chung et al., 2022) and mT5 (Xue et al., 2021) models in the sizes Small, Base, XL and XXL, allowing us to study effects over different model scales. We selected the T5 family of models over other models for three reasons: 1) Flan-T5 XL and XXL perform best in their parameter range on LegalBench, 2) T5 and mT5 allow us to measure the effect of multilinguality in a controlled setting, and 3) the T5 model family contains models from 60M parameters (Small) to 11B (XXL) allowing us to study scaling behaviour also at smaller scales.
@@ 2.1 Continued Pretraining
We continue pretraining on the MultiLegalPile (Niklaus et al., 2024), a 689GB corpus in 24 languages from 17 jurisdictions. It includes diverse legal data sources with varying licenses and allows for pretraining NLP models under fair use, with more permissive licenses for the Eurlex Resources and Legal mC4 subsets. It consists of four large subsets: a) Native Multi Legal Pile (112 GB), b) Eurlex Resources (179 GB), c) Legal mC4 (106 GB), and d) Pile of Law (292 GB). For our mT5 experiments, we use the entire corpus, and for T5 and Flan-T5 experiments, we use only English texts.
We continued pretraining (a.k.a. domain adaptation of) with 512 tokens in both inputs and targets on the MultiLegalPile (Niklaus et al., 2024) whereas the original models were pretrained on C4 (Raffel et al., 2020). We used the UL2 mixture (Tay et al., 2022) due to its promise to enable improved training efficiency with its mixture of denoisers. In initial experiments we used batch size 1024 and warmed up the learning rate linearly for the first 10K steps from 2.5e-3 to 5e-3, then decayed it to 1.5e-3. However, we noticed training instabilities for the XXL models. We switched to a constant learning rate of 1e-3 and ran a sweep over batch sizes 64, 128, 256, 512, 1024. The XXL model trained stably only with batch size 128.
@@ 2.3 Evaluation
We evaluate our models on LegalBench and MMLU to test in-domain and generalization performance, respectively. LegalBench (Guha et al., 2023) consists of 162 tasks evaluating different aspects of legal classification and reasoning. Each task is assigned to one of five categories, depending on the broader type of legal reasoning implicated. LegalBench tasks are sourced from both previously constructed datasets and novel tasks collected from different members of the legal community (e.g., lawyers, legal impact organizations, legal academics). As such, LegalBench is thought to capture tasks of interest and practical applicability. LegalBench tasks span a wide range of legal subject areas (e.g., contracts, civil procedure, tax, etc.) and text-types (natural language, contractual terms, judicial opinions, etc.). The majority of tasks are either classification or extraction tasks, thus enabling automated evaluation. Massively Multilingual Language Understanding (MMLU) benchmarks models factual knowledge (Hendrycks et al., 2021b). MMLU contains multiple-choice questions on 57 subjects, including three related to law: jurisprudence, international law, and professional law. We conducted preliminary experiments on the multilingual LEXTREME benchmark (Niklaus et al., 2023) but encountered significant obstacles: a) Many tasks in LEXTREME are NER-based, posing evaluation challenges for generative models like ours and b) The extreme multi-label classification tasks in LEXTREME showed near-random performance without task-specific fine-tuning, which is beyond the scope of our current work. Therefore, we focus on LegalBench and MMLU, both in English.
For evaluation, we set temperature to 0 in line with accepted practice for LegalBench evaluation (Guha et al., 2023) that focuses on the highest-likelihood token sequence with minimal variance. We removed the following prefixes before scoring: “label”, “target”, “option”, “answer”, “a:”. We did not evaluate Rule QA because it necessitated manual evaluation. We show paper baseline results compared with our runs in Appendix E Table 5. Our XL model is quite close to the XL model in the LegalBench paper, but there are significant differences for the XXL model. We provide a more detailed analysis of possible causes in Section C.1. Unless specifically mentioned, we compare to our baselines results. We hold out LegalBench tasks overlapping with LawInstruct tasks unless specified otherwise (see Section C.2 for details).
@@ 3 Results
This section discusses the main results from instruction tuning and continued pretraining Flan-T5.
@@
Figure 5: Performance progression on LegalBench and MMLU from baseline to instruction tuning (IFT) and continued pretraining followed by instruction tuning (PRE-IFT+IFT).
Table 1: Progression of performance from baseline to instruction tuning (IFT) and continued pretraining followed by instruction tuning (PRE-IFT+IFT).
@@
To analyze the change in performance in more detail, we show the difference to the baseline for the XL model on LegalBench and MMLU across tasks (see Figure 6) and across categories (see Figure 7). We find that FLawN-T5 outperforms baseline Flan-T5 in most LegalBench tasks in most categories. The exception are tasks in the interpretation category, specifically CUAD (Hendrycks et al., 2021c), where the fine-tuned model is actually worse than the baseline by around 10 points on average. A possible explanation could be negative transfer from the instruction tuning data since the task formulations are very different to the instructions in LegalBench. In MAUD (Wang et al., 2023a) and Contract-NLI (Koreeda and Manning, 2021), the instructions are much more similar from LawInstruct to LegalBench, leading to improvements compared to the baseline. On MMLU, most categories and tasks see increases in performance, especially the categories social sciences and other. We find that performance suffers mostly in the STEM category and to some extent in the humanities. Interestingly, the largest drop is in machine learning but the largest rise is in high school computer science. In the humanities, more “hard” disciplines are affected by performance decrease, such as formal logic and logical fallacies.
@@ 4 Ablations
In this section, we perform controlled experiments across the starting checkpoints, data mixtures, instruction styles and amount of instruction tuning data during pretraining. We show additional ablations regarding sampling styles, licenses and crosslingual transfer from multilingual data in Appendix D. Flan-T5 performs best in the studied parameter ranges. Baselines for other models are in Appendix E Table 5.
@@ 4.1 Starting Checkpoint
Should you start in-domain instruction tuning from a base model or from an instruction tuned model? $\Rightarrow$ Starting from an instruction tuned model is better across sizes except Small. In Figure 8, we compare instruction tuning from a base T5 and a Flan-T5 model in four different sizes (Small, Base, XL and XXL) (detailed results in Appendix E Table 6). We find that for the larger sizes, the instruction tuned Flan-T5 is a better starting point ($p<0.001$), leading to higher performance on LegalBench. For the Small size the difference is not statistically significant ($p=0.058$). We use the Flan-T5 model as a starting point in all experiments unless specified otherwise.
@@ 4.2 Data Mixture
What data mixtures should you choose for in-domain instruction tuning? $\Rightarrow$ Mixing in general instruction tuning datasets is necessary. In Figure 9, we compare instruction tuning with three different data mixtures: lawinstruct, flan2 (Chung et al., 2022), and flan2-lawinstruct (where we sample equally from flan2 and lawinstruct) (detailed results in Appendix E Table 7). Interestingly, when only training on lawinstruct, downstream accuracy drops, possibly due to the instructions in our datasets being formulated differently than the original Flan instructions. Training on flan2 and flan2-lawinstruct leads to an aggregate increase of 7.7 points (48.3 to 56) and 10.8 points (48.3 to 59.1) respectively. We use the flan2-lawinstruct mixture in all experiments unless specified otherwise.
@@ 4.3 Instruction Style
Are models trained with more diverse instructions better on LegalBench? $\Rightarrow$ Results are mixed, overall just using one instruction is probably sufficient. In Figure 10, we compare the performance of training with just one manually written instruction vs. ten paraphrased instructions with GPT-4 from one seed instruction, all else constant (detailed results in Appendix E Table 10). For Flan-T5 (see Table 10), for Small, one instruction is better than ten ($p=0.035$); for the other sizes we find no difference. For mT5 (see Figure 10(b)), for Small, one instruction is worse than ten both monolingual ($p=0.005$) and multilingual ($p=0.01$) whereas for XL, ten English instructions underperform one English ($p<0.001$) and ten multilingual ones ($p<0.001$). In aggregate, differences are small without a consistent trend.
@@ 4.4 Amount of Instruction Data During Continued Pretraining
How much instruction tuning data should be mixed in during continued pretraining? $\Rightarrow$ Continued pretraining seems to be rather robust w.r.t. the amount of instruction tuning samples mixed in. In Tables 12, 13, 14 and 15, we investigate the benefit of mixing varying amounts of instruction tuning data in during continued pretraining (detailed results in Appendix E Tables 12, 13, 14 and 15). We compare results on LegalBench of instruction tuning runs after 10K to 90K steps of continued pretraining. For the Small model, the benefit of continued pretraining over just instruction tuning is significant (34.9 for just instruction tuning vs.  40 after continued pretraining). Conversely, for the XL model, continued pretraining often underperforms compared to just instruction tuning. For the XXL model, more instruction tuning samples during continued pretraining improve performance, unlike for the Small and XL models. Across sizes, continued pretraining’s effectiveness appears robust to the number of instruction tuning samples used. [^4] Mixing instruction tuning data during continued pretraining without more instruction tuning does not improve results.